### v2

# Unraveling Syntax: Language Modeling and the Substructure of Grammars

Abstract
While large models achieve impressive results, their learning dynamics are far from understood. Many domains of interest – such as natural language syntax, coding languages, arithmetic problems – are captured by context-free grammars (CFGs). In this work, we extend prior work on neural language modeling of CFGs in a novel direction: how language modeling behaves with respect to CFG substructure, namely “subgrammars". We first define subgrammars, and prove a set of fundamental theorems regarding language modeling and subgrammars. We show that language modeling loss (or equivalently the Kullback-Leibler divergence) recurses linearly over its top-level subgrammars; applied recursively, the loss decomposes into losses for “irreducible" subgrammars. We also prove that the constant in this linear recurrence is a function of the expected recursion, a notion we introduce. We show that under additional assumptions, parametrized models learn subgrammars in parallel. Empirically, we confirm that small transformers learn subgrammars in parallel, unlike children, who first master simple substructures. We also briefly explore several other questions regarding subgrammars. We find that subgrammar pretraining can improve final performance, but only for tiny models relative to the grammar, while alignment analyses show that pretraining consistently lead to internal representations that better reflect the grammar’s substructure in all cases; we also observe persistent difficulty with deeper recursion, a limitation that appears even of large language models.
Context-free grammars, Probabilistic context-free grammars, transformer, subgrammar, learning dynamics, formal language, representation space
Equal contribution
1ETH Zürich
2Massachusetts Institute of Technology
1 Introduction
Large language models (LLMs) have stunned the world by achieving sophisticated language abilities in the past few years, yet we still do not know how they reach such high levels of performance. Little is also known about the process of language acquisition. Do LLMs, for example, master simpler substructures before progressing to more complex syntax, as children do?
A major approach has been to study trained language models (for instance investigating how a trained model analyzes and uses its knowledge of a language during inference). More recently, a small but burgeoning direction has studied how neural architectures learn Context-Free Grammars (CFGs), a class of formal languages that broadly captures most domains of interest, such as natural languages and programming languages. The key insight is that by training models on smaller, fully controllable CFGs, training can be efficient, and one can probe for features of CFGs (specific rules, grammar size, etc). These approaches have gained us many valuable insights (as discussed in the Related Work).
However, two things have been largely unstudied until now. First, little has been shown about the dynamics of how models acquire language – not the static representations or logic of trained models (even in the CFG / synthetic language literature). Second, research studying CFGs has not considered that CFGs as a mathematical object have substructure; they decompose into “subgrammars". Indeed, in analogous research areas that study learning of abstract hypothesis classes such as polynomials, XOR functions, and modular counting, a major focus has been studying how learning interacts with the substructure of these function classes (e.g. the monomials that compose polynomials).
In this work, we advance the study of language modeling of CFGs by analyzing it through the subgrammar structure. Many of our results can also be viewed through the lens of studying the dynamics of CFG learning. In Section 4, we begin by defining several notions of subgrammars: inner subgrammars, corresponding to subtrees of CFG derivations, and outer subgrammars, corresponding to simplified versions of the CFG. Our definitions of subgrammars in this way are novel (though related to classic work on the algebra of CFGs), and we believe they are the right notions for studying the substructure of CFGs. The most important contribution of our work is a suite of fundamental theorems showing that the loss of language modeling, or equivalently the Kulback-Leibler (KL) divergence), obeys a recurrence over the subgrammar structure. Under additional (strong) assumptions this implies a model learns subgrammars in parallel. Empirically, we show that small transformers trained on CFGs learn all the subgrammars in parallel, unlike how children acquire language.
Next, changing gears in 5 we study whether curriculum learning, by using an inductive bias and training on a subgrammar first, can improve performance: for small models, we show it can. In 5.2 we use alignment analysis to show, quite definitively, that such pre-training results in very different internal representations of the CFG: it aligns subgrammar strings, and non-subgrammar strings, respectively, resulting in internal representations that reflect the substructure of the CFG. Finally, in Section 6 we show expeirmentally that even models that perform well do not “know” the subgrammar structure perfectly, with the depth of recursion being the main difficulty.
2 Related Work
Transformers (Vaswani et al., 2017), and language models more broadly, have been studied in two predominant research directions: improving training methods (Bubeck et al., 2023; Jaech et al., 2024; Guo et al., 2025) and probing trained models to analyze how knowledge is stored and activated during inference (Meng et al., 2022; Geva et al., 2021; Dar et al., 2022; Ferrando and Voita, 2024). Much less is known about how such models acquire language, though evidence of stage-like acquisition reminiscent of child language learning in GPT-2 has been reported Evanson et al. (2023).
We approach this problem via the surrogate (and theoretically significant in its own right) approach of studying the dynamics of language models acquiring formal languages. This complements prior work on gradiient-based learing over structured hypothesis classes (e.g. juntas, parities, modular counting) (Klivans and Kothari, 2014; Telgarsky, 2016; Abbe et al., 2024; Daniely and Malach, 2020). CFGs provide a linguistically motivated setting where recursive structure is explicit, and formal language theory offers a well-developed foundation (Cotterell et al., 2023).
Formal languages have been used to test neural models, with mixed success. RNNs and LSTMs often fail to learn subregular grammars despite theoretical capacity (Avcu et al., 2017), and transformers perform well on many formal languages but struggle with recursion and counter-based mechanisms (Bhattamishra et al., 2020). Other studies confirm that transformers often fail on deeply nested grammatical structures (Lampinen, 2024). Results consistently show that gradient descent, rather than model expressivity, is the limiting factor. Theoretically, self-attention has limitations for some long-range dependencies (Hahn, 2020) despite transformers’ expressivity results (Pérez et al., 2021; Yun et al., 2019) (see Strobl et al., 2024 for a survery).
Closest to our focus, (Cagnetta and Wyart, 2024) study transformers trained on PCFG-generated data and relate learning behavior to hierarchical structure in the underlying grammar, while (Allen-Zhu and Li, 2023) provide a thorough analysis of how trained transformers can implement CFG-like computations and encode rewrite rules in their internal states. Our work builds on this by studying the learning dynamics, specificall vis-a-vis subgrammars.
3 Preliminaries and Definitions
3.1 Formal Languages
Definition 3.1 (CFG).
A Context-Free Grammar (CFG) is a tuple $G=(\Sigma,\mathcal{N},\mathcal{S},\mathcal{P})$ where $\Sigma$ is a finite set of terminal symbols, $\mathcal{N}$ is a finite set of non-terminal symbols, $S\in\mathcal{N}$ is the designated start symbol, and $\mathcal{P}$ is a finite set of production rules of the form $A\rightarrow\alpha$ where $A\in\mathcal{N}$ and $\alpha\in(\mathcal{N}\cup\Sigma)^{*}$ ($\alpha$ can be the empty string $\epsilon$).
The language $L_{G}\subseteq\Sigma^{*}$ associated with a CFG $G$ is the set of all strings of terminals derivabl from $S$ using rules in $\mathcal{P}$.
Definition 3.2 (PCFG).
A Probabilistic Context-Free Grammar (PCFG) is a context-free grammar $G=(\Sigma,\mathcal{N},\mathcal{S},\mathcal{P})$ augmented with a probability function $\mathcal{W}$ that assigns to each rule $(A\rightarrow\alpha)\in\mathcal{P}$ a probability, such that for each $A\in\mathcal{N}$, $\sum_{\{(A\rightarrow\alpha)~\in~\mathcal{P}\}}\mathcal{W}(A\rightarrow\alpha)=1$.
CFGs were originally defined in the context of linguistics (Chomsky, 1956), as the vast majority of the syntax of natural languages, as well as the syntax of programming languages and mathematics, can be formulated as CFGs (Shieber, 1985; Pullum and Gazdar, 1982). Since CFGs capture languages with recursion and embedded structure, there intuitively exists a notion of a subgrammar. However, several subtleties crop up when attempting to define subgrammars. We propose two notions of subgrammars, each of independent interest and relevance: one is the grammar of substrings of CFG sentences that can be generated from a non-terminal, and the other as a subset of the CFG language generated by a subset of the rules. We term these inner and outer subgrammars respectively. Intuitively, inner subgrammars correspond to subtrees of derivations of CFGs, whereas outer subgrammars are a simplified version of the grammar. We will sometimes say supergrammar for a bigger grammar containing a subgrammar.
Definition 3.3 (Inner Subgrammar).
An inner subgrammar of a PCFG $G=(\Sigma,\mathcal{N},\mathcal{S},\mathcal{P},\mathcal{W})$ is itself a PCFG $G^{\prime}=(\Sigma^{\prime},\mathcal{N}^{\prime},\mathcal{S}^{\prime},\mathcal{P}^{\prime},\mathcal{W}^{\prime})$ such that $\Sigma^{\prime}\subseteq\Sigma$, $\mathcal{N}^{\prime}\subseteq\mathcal{N}$, and $\mathcal{P}^{\prime}$ is the set of all rules expanding non-terminals in $\mathcal{N}^{\prime}$. Finally $\mathcal{W}^{\prime}$ is the restriction of $\mathcal{W}$ to $\mathcal{P}^{\prime}$, i.e. renormalized so that for every $A\in\mathcal{N}^{\prime}$, $\sum_{\{(A\rightarrow\alpha)\in\mathcal{P}^{\prime}\}}\mathcal{W}^{\prime}(A\rightarrow\alpha)=1$ .
Definition 3.4 (Proper Subgrammar).
A proper subgrammar is an inner subgrammar $G^{\prime}$ of a CFG $G$ which which is not the whole grammar (in particular $S\notin\mathcal{N}^{\prime}$).
Definition 3.5 (Outer Subgrammar).
An outer subgrammar of a PCFG $(\Sigma,\mathcal{N},\mathcal{S},\mathcal{P},\mathcal{W})$ is a PCFG $G^{\prime}=(\Sigma^{\prime},\mathcal{N}^{\prime},\mathcal{S},\mathcal{P}^{\prime},\mathcal{W}^{\prime})$, with $\Sigma^{\prime}\subseteq\Sigma$, $\mathcal{N}^{\prime}\subseteq\mathcal{N}$, $\mathcal{P}^{\prime}\subseteq\mathcal{P}$, where $\mathcal{W}^{\prime}$ is the renormalized restriction of $\mathcal{W}$ to $\mathcal{P}^{\prime}$. To be a valid outer subgrammar, $\mathcal{P^{\prime}}$ must contain at least one rule from $\mathcal{P}$ where the left-hand side is $S$.
An outer subgrammar captures the notion of a subset of the language generated by a PCFG obtained by keeping a subset of expansions of some non-terminals (including $S$). Every string generated by an outer subgrammar is a valid string of the supergrammar. An outer subgrammar more closely corresponds to the notion of a “simple" version of a language, whereas inner subgrammars are the inherent compositional substructures of a CFG.
3.2 Language Modeling
In this work, all distributions are assumed to be over strings of a finite alphabet $\Sigma$, although most definitions can be extended to arbitrary domains.
Definition 3.6 (Kullback-Leibler Divergence).
Given distributions $P$ and $Q$ over $\Sigma^{*}$, the Kullback-Leibler (KL) Divergence of $Q$ from $P$ is
$\mathrm{KL}(P\;\|\;Q)=\sum_{s\in\Sigma^{*}}P(s)\log\frac{P(s)}{Q(s)}$
A language model $Q_{\theta}$ is a function family parametrized by $\theta$, such that $Q_{\theta}(x)$ yields a probability distribution over $x\in\Sigma^{*}$. In this work one can think of all $Q_{\theta}$ as auto-regressive (though for several theoretical results this is not strictly necessary), meaning $Q_{\theta}(x)$ is an abstraction of a next-token prediction model, i.e. $Q_{\theta}(x_{1},\ldots,x_{n})=\Pi_{i=1}^{n}Q_{\theta}(x_{i}|x_{1},\ldots,x_{i-1})$.
In Language Modeling, $Q_{\theta}$ is optimized with Maximum Likelihood Estimation:
Definition 3.7 (Maximum Likelihood Estimation).
Given a target distribution $P$, the Maximum Likelihood Estimator $Q_{\hat{\theta}}$ is parametrized by $\hat{\theta}=\text{arg max}_{\theta}\mathcal{L}(\theta)$ where
$\mathcal{L}(\theta)=\mathbb{E}_{s\sim P}~[-\log Q_{\theta}(s)]$
Practically, this is done by maximizing the combined likelihood of a set of samples, or equivalently, minimizing the sum of negative log-likelihoods; in the limit, this exactly approaches $\mathcal{L}(\theta)$.
Definition 3.8 (Shannon Entropy).
The Shannon Entropy of a probability distribution is
$H(P)=\mathbb{E}_{s\sim P}~[\log P(s)]$
Proposition 3.9.
Given a true distribution $P$ and model $Q_{\theta}$ parametrized by $\theta$,
$\mathcal{L}(\theta)=D_{\mathrm{KL}}(P~\|~Q_{\theta})+H(P)$
The proof (given in the Appendix B) is a straightforward application of the linearity of expectation. In particular, this implies that $\hat{\theta}$ minimizes $\mathcal{\theta}$ if and only if it minimizes $D_{\mathrm{KL}}(P~\|~Q_{\theta})$.
4 The Fundamental Relation of Language Modeling and Subgrammars
Theorem 4.1 (Unique decomposition).
Every (P)CFG $G$ can be uniquely decomposed into a hierarchy of its inner subgrammars.
This hierarchical structure can be represented as a directed acyclic graph (DAG) with self-loops, where each node is labeled by the set of non-terminals that generate the corresponding subgrammar.
The straightforrward proof recursively constructs the DAG by first identifying the “top-level" subgrammars of $G$; see Appendix B. While to our knowledge this theorem in its formulation is our own, the nodes of the DAG decomposition correspond to the “grammatical levels" of a CFG in Gruska’s classical work on CFG theory (Gruska, 1971).
4.1 Subgrammars and Language Modeling
We now study the connection between the subgrammar structure and language modeling. Let $G=(\Sigma,\mathcal{N},S,\mathcal{P},\mathcal{W})$ be a PCFG that induces a distribution $P_{G}$ over $\Sigma^{*}$, and $Q_{\theta}$ an autoregressive language model trained to approximate $P_{G}$ (that is, given a sequence in $\Sigma^{*}$ it outputs a terminal, or EOS).
For motivation, first consider the simple case where the only expansion of $S$ is $S\rightarrow\alpha A\beta$, where $A$ is some proper subgrammar (does not generate $S$), and $\alpha,\beta\in\Sigma^{*}$ are strings of terminals. Then,
$D_{\mathrm{KL}}(P~\|~Q)=\sum_{a\in\Sigma^{*}}P_{G}(\alpha a\beta)\log\frac{P_{G}(\alpha a\beta)}{Q_{\theta}(\alpha a\beta)}$
$\frac{\log P_{G}(\alpha|\epsilon)}{\log Q_{\theta}(\alpha|\epsilon)}+\sum_{a}P_{G}(a)\frac{\log P_{G}(a)}{\log Q_{\theta}(a|\alpha)}\\ +\sum_{a}P_{G}(a)\frac{\log P_{G}(\beta|\alpha a)}{\log Q_{\theta}(\beta|\alpha a)}$
In an abuse of notation, above $P_{G}(\alpha|\epsilon)$ denotes the probability of a (partial) sequence that starts with $\alpha$, $P_{G}(a|\alpha)$ the probability of $a$ following $\alpha$, and so on. Importantly, the decomposition of $P_{G}$ and $Q_{\theta}$ follows from the subgrammar structure of $G$ in the case of $P_{G}$, and from the fact that $Q_{\theta}$ is autoregressive for the $Q_{\theta}$ terms. In short, the KL-divergence evaluates to a sum of “conditioned KL-divergences" of the subgrammar $A$, of prefix $\alpha$, and of suffix $\beta$. The latter terms can themselves be thought of as simple subgrammars; indeed, we can rewrite $G$ to include two new non-terminals that evaluate to $\alpha$ and $\beta$ respectively (with prob. 1), and we would then have a sum over three “sub-divergences".
Definition 4.2.
Given PCFG distribution $P_{G}$, arbitrary distribution $Q$ over $\Sigma^{*}$, and top-level subgrammar $A$ of $G$, define $D_{\mathrm{KL}}(P_{G}~\|~Q)_{A}$
$\sum_{s\in\Sigma^{*}}P(s|\epsilon)P_{G}(A|s)\sum_{a\in\Sigma^{*}}D_{\mathrm{KL}}(P_{G}~\|~Q(\cdot|s))$
$D_{\mathrm{KL}}(R~\|~Q)_{A}$ can be seen as the “restriction" of the KL-divergence to the subgrammar $A$ (by averaging over all contexts that can begin $A$). In the case of a fixed string $\alpha\in\Sigma^{*}$ we will write $D_{\mathrm{KL}}(P_{G}~\|~Q)_{\alpha}$ where the second sum is replaced with a single term for $\alpha$ (as one can view $\alpha$ as a subgrammar of one string). Then we have, in our previous example, $D_{\mathrm{KL}}(P_{G}~\|~Q)=D_{\mathrm{KL}}(P_{G}~\|~Q)_{\alpha}+D_{\mathrm{KL}}(P_{G}~\|~Q)_{A}+D_{\mathrm{KL}}(P_{G}~\|~Q)_{\beta}$.
Our first fundamental result is that this decomposition, or recurrence, holds generally. Let the top-level subgrammars denote the children of the root node in a CFG’s subgrammar decomposition.
Theorem 4.3 (KL loss as a recursive function over subgrammars).
Let $G$ be a PCFG with top-level subgrammars $A_{1},\ldots,A_{k}$. Let $C\subset\Sigma^{*}$ be the set of (fixed) substrings of terminals in expansions of $S$. Then $D_{\mathrm{KL}}(P_{G}~\|~Q_{\theta})$
$\sum_{i=1}^{k}D_{\mathrm{KL}}(P_{G}~\|~Q_{\theta})_{A_{i}}~+\sum_{\alpha\in C}D_{\mathrm{KL}}(P_{G}~\|~Q_{\theta})_{\alpha}$
Corollary 4.4.
If we rewrite $G$ as an equivalent PCFG with additional non-terminals such that $S$ maps to strings only non-terminals (corresponding to subgrammars $A_{1},\ldots,A_{k})$; then the right sum of Theorem 4.3 can be removed:
$D_{\mathrm{KL}}(P_{G}~\|~Q_{\theta})=\sum_{i=1}^{k}D_{\mathrm{KL}}(P_{G}~\|~Q_{\theta})_{A_{i}}$
The full proof of Theorem 4.3 and Corollary 4.4 are in Appendix B. Upon closer inspection, it is clear that the recursion actually applies to any subgrammar; that is, for subgrammar $A$ with subgrammars $B_{1},\ldots,B_{l}$, $D_{\mathrm{KL}}(P_{G}~\|~Q_{\theta})_{A}=\sum_{j}D_{\mathrm{KL}}(P_{G}~\|~Q_{\theta})_{B_{j}}$ (indeed, we could have states Theorem 4.3 with respect to subgrammars, as $D_{\mathrm{KL}}(P_{G}~\|~Q_{\theta})=D_{\mathrm{KL}}(P_{G}~\|~Q_{\theta})_{G}$). Hence, this formula can be expanded recursively over each of the subgrammars $A_{i}$ by repeated application, resulting in a sum over all the leaves of the DAG decomposition of $G$ into its subgrammars; see Corollary B.1 in the Appendix for the precise statement.
Now, suppose each top-level subgrammar $A_{i}$ occurs with probability $p_{i}$ over the top-level rules that expand $S$; it is tempting to conclude that the recursive formula simplifies to $KL(P_{G}~\|~Q_{\theta})=\sum_{i=1}^{k}p_{i}D_{\mathrm{KL}}(P_{G}~\|~Q_{\theta})$ (where the KL terms are no longer restrictions, but bona-fide divergences between the distribution $P_{G}$ and $Q_{\theta}$ as a language model for $A$). However, this is true only if $Q_{\theta}$ is excellent and models $P_{A}$ identically under any context where the subgrammar $A$ can occur, which may not be the case!
Corollary 4.5.
Let $G$ be a PCFG where $S$ evaluates to rules with only non-terminals (correspondingly, subgrammars) $A_{1},\ldots,A_{k}$ each of which occurs with prob. $p_{i}$.
Assume $Q_{\theta}$ is “context insensitive" for each grammar $A_{i}$: that is, for two contexts $s,s^{\prime}$ for which $P_{G}(A_{i}|s)P_{G}(A_{i}|s^{\prime})>0$, $Q_{\theta}(A_{i}|s)=Q_{\theta}(A_{i}|s^{\prime})$ (the restrictions of $Q_{\theta}$ to strings from $A_{i}$ given possible contexts $s$ or $s^{\prime}$, are the same). Then
$D_{\mathrm{KL}}(P_{G}~\|~Q_{\theta})=\sum_{i=1}^{k}p_{i}D_{\mathrm{KL}}(P_{A_{i}}~\|~Q_{\theta}(A_{i}))$
Where $Q_{\theta}(A_{i})=Q_{\theta}(A_{i}|s)$ for arbitrary context $s$ s.t. $P_{G}(A_{i}|s)>0$.
Several comments are in order about this Corollary, which simplifies the general recursive formula of Theorem 4.3 so that the recursive terms are simple KL-divergences (not “conditioned" KL-divergences). The corollary requires that the strong assumption that the model be "context insensitive" for its subgrammars, but results in a particularly elegant decomposition. When one considers a model being trained over time, it may or may not context insensitive for a given subgrammar at different steps; but at any point that it is, this fundamental recurrence must hold. While we do not present it formally out of interest of space, it is not hard to extend to approximate / statistical versions of this corollary: to the extent that $Q_{\theta}$ is not context-insensitive, the difference between the elegant decomposition and the true loss will differ to the same extent. Finally, our experiments suggest that this condition is perhaps not so strong, at least in the statistical sense: in the experiments in Figures 1, discussed below, qualitatively similar results were obtained when we computed subgrammar divergences with varying prefixes. In Section 6, we find that for prefixes of increasing length, our small transformer models the distribution of the ensuing subgrammar identically, but not if the prefixes are highly deep; however, such strings are “rare" under the actual distribution (so one could indeed say that these models appear to be “context-insensitive statistically").
In another direction, consider that in Theorem 4.3 and its corollaries, any of the top-level subgrammars $A_{i}$ could have been the grammar $G$ itself (if $G$ has a self-loop). It turns out we can say even more about the KL-divergence as a function of the degree of “self-loopiness", or recursion.
Theorem 4.6 (KL-divergence with expected recurrence).
Let $G$ have proper top-level subgrammars $A_{1},\ldots,A_{k}$, each occurring with prob. $p_{k}$ over rules expanding $S$, and let $Q_{\theta}$ be context-insensitive for subgrammars of $G$.
Let the recursion R be the number of times $S$ occurs in the rule that expands $S$. Then,
$D_{\mathrm{KL}}(P_{G}~\|~Q_{\theta})=\frac{\sum_{i=1}^{k}D_{\mathrm{KL}}(P_{A_{i}}~\|~Q_{\theta}(A_{i}))}{1-\mathbb{E}[R]}$
If $1-\mathbb{E}[R]<0$, then the KL-divergence is unbounded if $D_{\mathrm{KL}}(P_{A_{i}}~\|~Q_{\theta}(A_{i}))>0$ for any $A_{i}$.
See B for the full proof. Theorem 4.6 can be seen as the equation for the “base case" in the recursive formula for KL-divergence, since an irreducible (leaf) subgrammar evaluates only to strings of terminals and itself! This equation shows that the expected recursion in such a (sub)grammar must be less than 1 (and the closer it is to 1, the greater the “blow-up" of its divergence to a language model); indeed, if the expected recursion is 1 or greater, the PCFG sampling process that recursively expands the root symbol will in expectation never terminate. Note that a similar, but more clunky, theorem can be shown without assuming context-insensitivity to subgrammars (replacing KL-divergences with conditioned / averaged versions).
Finally, Theorem B.2 proves a similar, recursive decomposition for outer subgrammars; the statement and proof have been moved there for brevity.
4.2 Experiments and Visualizations
To visualize these theorems, we train small transformers on synthetic CFGs with varied subgrammar structure, and plot the KL-divergence over training in Figure 1. This example plot shows visually how, throughout all stages of learning, the KL divergence (loss) is the sum over the corresponding loss for each subgrammar.
Figure 1: A grammar with inner subgrammars, each occurs with 100% probability. Overhead refers to constant strings in between subgrammar roots.
(a) Deeper Recursion: a language with an inner subgrammar DAG of depth 4.
(b) KL decomposition for an outer subgrammar using most of the rules (see Theorem B.2)
Figure 2: Additional examples of how loss, or KL-divergence, behaves with respect to varying subgrammar structure.
To illustrate Theorem 4.6, consider a simple CFG with two rules:
$S\rightarrow x\;\;(p),\quad S\rightarrow(S~\text{and}~S)\;\;(1-p)$
The expected recursion is $\mathbb{E}[R]=2(1-p)$. Assuming the language modelis context-insensitive, we then the KL-divergence is C/(2p1)C/(2p-1) for some constant $C$. We train a small transformer over this language with increasing $p\in(0.5,1]$, demonstrating qualitatively the non-linear (inverse proportional) growth of KL-divergence as $p$ (the prob. of not recursing) approaches $0.5$; see Figure 4. Finally, we use similar experiments to visualize loss decomposition for a grammar with many nested subgrammars (i.e. the DAG is a line), and for outer subgrammars, in Figure 2.
However, an additional phenomenon jumps out from all of these plots: the models learn all subgrammars in parallel! One might have intuitively expected a model to first master a simpler subgrammar before progressing to the encompassing supergrammar. While the loss decomposition theorems show that nothing within the task of language learning prevents parallel learning of subgrammars, it is possible to construct pathological, theoretical scenarios where a model independently optimizes subgrammars in sequence. Parallel subgrammar learning is property of the training method and model architecture. We believe our work opens a fascinating new direction of studying when and why models learn all subgrammars in parallel. Here, we present a simplistic but fundamental scenario in which this occurs:
Corollary 4.7.
(Stated informally) Suppose $Q_{\theta}$ is trained on PCFG $G$ with subgrammars $A_{1},\ldots,A_{k}$ via gradient descent, and that the model and PCFG together obey a kind of “independence": for a gradient update to $\theta$ on a subgrammar $A_{i}$, that is $\delta=\nabla_{\theta}(-D_{\mathrm{KL}}(P_{G}~||~Q_{\theta})_{A_{i}})$, applying it does not hinder performance on other subgrammars. That is, for $\theta^{\prime}=\theta+\delta$, $D_{\mathrm{KL}}(P_{G}~||~Q_{\theta^{\prime}})_{A_{j}}\leq D_{\mathrm{KL}}(P_{G}~||~Q_{\theta})_{A_{j}}$ for $j\neq i$ (in fact it is sufficient for this condition to hold only for $\theta$ within the path of descent).
Then, if trained with gradient descent, $Q_{\theta}$ learns all subgrammars in parallel.
Note that this is indeed a Corollary; it would not always be true if loss did not recurse linearly over subgrammars. one immediate future direction would be to study whether the small transformers and PCFGs of this paper learn subgrammars in parallel because they satisfy the independence condition of 4.7; we believe this to be the case. Despite their small size, they are likely still overparametrized with respect to the even tinier PCFGs. Future work can also aim to weaken the assumptions for parallel learning.
5 Subgrammars and Improving Optimization
While the previous section establishes a mathematical relationship between training loss and subgrammar structure, it is natural to consider whether the structure of CFGs could be exploited in training; e.g. is pretraining on a subgrammar helpful? Perhaps mastering simpler components first facilitates learning of more complex structures later. Such approaches are studied in curriculum learning (Bengio et al., 2009; Wang et al., 2021) and modular pretraining strategies (Andreas et al., 2016; Kaiser et al., 2017).
Two-layer Transformer Four-layer Transformer
Pretraining 10 epochs Pretraining 20 epochs Pretraining 10 epochs
Attention MLP Attention MLP Attention MLP
Full grammar sequences
From Scratch 0.258 0.535 0.249 0.535 0.249 0.469
With Pretraining 0.281 0.534 0.303 0.511 0.323 0.491
Percentage change (%) +8.9 -0.2 +21.7 -4.7 +8.3 +1.0
Subgrammar sequences
From Scratch 0.298 0.561 0.288 0.558 0.295 0.513
With Pretraining 0.339 0.566 0.348 0.544 0.347 0.525
Percentage change (%) +13.8 -0.1 +20.8 -2.6 +10.7 +1.9
Subgrammar pretraining only 0.288 0.558 0.288 0.558 0.295 0.523
Table 1: Average CKA similarity (0–1) across attention and MLP layers of a different Transformers when pretraining for 10 vs. 20 epochs. The most significant differences are in teal.
5.1 Robustness to Subgrammar Location
One might expect the choice of subgrammar to influence learning, given the autoregressive nature of transformers. In particular, a prefix subgrammar, an inner subgrammar always occurring at the beginning of sequences of G, might be easier to retain, whereas the results from pretraining on a suffix subgrammar or an infix subgrammar (appearing in the middle and disconnected from sentence endpoints) might be overwritten when training on the full grammar begins. However, our results show this is not the case: the model reliably retains modeling performance on any subgrammar, regardless of its position. This robustness is illustrated in Figure 5. As the experiments of the following section suggest, it appears that training on a subgrammar ferries the model into a distinct area of weight space in which the subgrammar is internally represented, and further optimization (on the whole language) remains in this subspace.
5.2 Activation-space analysis
We examine how subgrammar pretraining affects internal representations by comparing models trained from scratch to those pretrained on a subgrammar and then continued on the full grammar. Similarity is measured with Centered Kernel Alignment (CKA) (Kornblith et al., 2019) across 30 random seeds.
Much to our surprise, we also found that for smaller models, subgrammar pretraining can even help achieve a lower final loss (Figure 6). This effect diminishes as the model size and representational complexity increase (for instance, this occurs for 2-layer transformers but not 4-layers). As expected, larger models consistently reach lower losses regardless of pretraining.
CKA analysis reveals that pretrained models exhibit higher alignment across attention layers than models trained from scratch, both when computed over full-grammar sequences, and (less surprisingly) subgrammar sequences (Table 1). A longer pretraining phase further increases alignment, although excessive pretraining can eventually reduce gains in final loss (see same Table).
Why are the pretrained models more “aligned" to one another (that is, represent sequences more similarly)? To probe this, we compare the representational similarity of the top quantile of seeds via cosine similarity of embeddings of three types of sequences: (i) sequences consisting solely of the subgrammar, (ii) sequences with no occurrence of the subgrammar, and (iii) sequences with both the subgrammar and other subsequences. We also compute (iv) the similarity between embedded pairs of a subgrammar sequence and a subgrammar-free sequence. For (i) and (ii), the attention-layers of pretrained models cluster subgrammar sequences (resp. no subgrammar sequences) significantly closer together than directly-trained models. This suggests that substructures learned during pretraining are retained after exposure to the full grammar. Finally, the gap between (iv) and (i), and between (iv) and (ii) is greater in pretrained models, suggesting pretrained models are better at internally segregating sequences with and without subgrammar subsequences (Table 3).
Our experiments are not exhaustive, and we leave open the question of how to train a model to consistently converge to the best optima, given the rather strong prior of the subgrammar structure of the target CFG. Too little pretraining may not provide a strong enough inductive bias, while too much may over-specialize the model to the subgrammar and hinder transfer. This trade-off mirrors classical insights from curriculum learning, where an optimal “window” of pretraining exposure exists (Bengio et al., 2009; Weinshall et al., 2018).
6 Generalization: Do LMs “Know Syntax"?
(a) Contexts of the form $(a)^{i}$ (depth 0)
(b) Contexts of the form $(^{i}$ (depth ii)
Figure 3: LM error vs. longer context, with or without recursion
With language models achieving low training loss, it is natural to ask whether they genuinely internalize and can generalize the rules of the PCFG. This question connects to the broader debate about whether language models exhibit intelligence in terms of structure and composition, or whether they are best understood as extraordinarily powerful pattern-matchers.
To probe this, we evaluate a small transformer trained on an especially simple PCFG: Nested Parentheses (Appendix C). The model achieves very low loss statistically. We test generalization to probabilistically unlikely (but grammatically valid) sequences with increasing length in two ways: (i) extending the context at the same depth of recursion, feeding in $(a)^{i}$, and (ii) growing sequences through repeatedly applying the recursive rule, resulting in contexts at increasingly deeper depths of recursion, of the form $)^{i}$. We then compare the model’s output logits (its output distribution) against the ground-truth next-token distribution. The next-token distribution is identical for all test contexts, even between cases (i) and (ii).
Figure 3 shows a striking contrast. For case (i), the prediction error remains low throughout, while for case (ii) it grows similarly to an inverse log curve. While the model appears to master the rules of the PCFG at shallow depth, this does not translate into robust handling of deeper recursive dependencies.
We also evaluate the effect of prepending different valid prefixes to the sequence of increasing depth. The results remain largely unchanged— even when using a faulty (non-grammatical) prefix. This suggests that the model’s primary difficulty lies in handling the depth of the subsequence it must complete, while it pays relatively little attention to the completed prefix.
Anecdotally, we find similar behavior even in state-of-the-art frontier models. We test GPT-5.1 Instant model on arithmetic expressions generated by a PCFG, presenting two kinds of long expressions: a chain composed of non-deep arithmetic operations, and a single deep arithmetic expression (depth 7) [^1] We do not find the same discrepancies for GPT-5.1 Thinking, which solves all of our examples within 3-4 minutes for each expression. The Thinking model may pass arithmetic expressions to a calculator or program, and/or uses an externally prompted or engineered chain-of-thought process; in any case, this departs from language modeling in the strict sense.. These experiments show  [^2] These arithmetic tests are purely anecdotal and should not be interpreted as direct evidence about training difficulty on recursive PCFGs. Our intention is to illustrate informally that deep recursion can stress current models, consistent with the difficulties observed in our experiments. that even LLMs, similar to our small LMs, struggle with depth and not length, correctly answering 5/5 non-deep arithmetic expressions but only 2/5 for a deep arithmetic expression. Note that for the not-deep arithmetic expressions (type 1), the LM in fact has to solve more terms than with the deeper recursion, but still solves them correctly.
7 Discussion and Future Work
Our work initiates the study of language-modeling dynamics with respect to the substructure of Context-Free Grammars. On the theoretical side we show that the language-modeling loss decomposes into subgrammar-local contributions. Empirically, we find that small transformers tend to reduce error across many subgrammars in parallel; we offer in a corollary a condition under which gradient descent training has this property, but more generally, we believe this phenomenon to be an exciting direction with open questions. We find subgrammar pretraining acts as an inductive bias: in tiny models it can improve performance, and even when it does not, it yields representations more aligned with the grammar’s substructure; understanding why, or whether this has any benefits is an open question. Finally, recursion experiments highlight depth – not length – as a persistent challenge in static (trained) language models.
8 Limitations
On the theoretical side, our results thoroughly characterize how language-modeling loss decomposes with respect to subgrammar structure in a variety of cases. However, we do not have a full theory for when or why gradient-based training yields parallel improvement across subgrammars; towards this, we present a single, somewhat heavy-handed result (Corollary 4.7. Empirically, our experiments use a small set of synthetic PCFGs and scaled-down, decoder-only transformers. Although these grammars are designed to isolate recursion, and subgrammar sizes / depth, and so on, they do not cover the full diversity of CFGs (in particular ambiguity – in all our experiments, CFGs are fully unambiguous; the effect of multiple possible parse-trees presents an interesting direction). Our results of Section 6 leave unresolved whether failures at deep recursion reflect representational limits or optimization barriers: we conjecture that there exists a setting of the weights of, say, a 2-layer, 2-head transformer (as in our experiments) that does correctly model the PCFG (at least up to some very high bound on depth). This would show that the issue is gradient descent which is not able to find such ideal solutions, analogous to work showing that while neural networks can in principle represent functions like parity, modular counting, or compositional rules, gradient descent often fails to find these solutions without strong inductive bias or curricula (Telgarsky, 2016; Abbe et al., 2024). Finally, we do not compare learning dynamics across other language classes in the Chomsky hierarchy (e.g., regular or mildly context-sensitive languages) under matched conditions, and thus cannot yet disentangle “difficulty of depth” in CFGs from other forms of dependency. Our work also does not explore the question of grammar induction, the learning task of determining the CFG underlying the input data.
Appendix A Details of our Transformer Architecture
The transformer architectures used in our experiments are scaled-down variants of nanoGPT (Karpathy, 2023). Training proceeds with batches sampled uniformly at random from the dataset. The number of batches per epoch depends on the total size of the training data – this implies that PCFG $G$ which generates longer sequences yield more iterations per epoch. Furthermore, the tokenizers contain only two special tokens: $BOS$ (beginning-of-sequence) and $EOS$ (end-of-sequence). We deliberately omit $UNK$ (unknown) and $PAD$ (padding) tokens, since all tokens are guaranteed to be in the grammar’s terminal set $\mathcal{N}$; this ensures the training distribution matches as closely as possible to the grammar distribution.
A.1 Model Parameter Settings
All models share the same decoder-only Transformer architecture as nanoGPT (Karpathy, 2023). Each model consists of a learned token embedding matrix $E\in\mathbb{R}^{|\mathcal{V}|\times d}$, learned positional embeddings for a fixed context window of $256$ tokens, $L$ stacked decoder blocks with multi-head self-attention and a two-layer feed-forward network with hidden size $4d$, followed by a final LayerNorm and a tied output projection $E^{\top}$. We use GELU activations, dropout rate $p=0.1$ in the attention, feed-forward, and embedding layers, and LayerNorm with learned scale and bias. Input and output token embeddings are tied, and we exclude the positional embeddings when reporting parameter counts.
We vary the number of layers $L\in\{1,2,4\}$, the model dimension $d\in\{6,8,20,32\}$, the number of attention heads $h\in\{1,2,4\}$ (with per-head dimension $d/h$), and the vocabulary size $|\mathcal{V}|$, which is determined by the underlying grammar. Table 2 summarizes the configurations used in our experiments.
Config $L$ $h$ $d$ $|\mathcal{V}|$
FourLayer 4 4 8 100
TwoLayer 2 2 20 100
TwoLayer_SMALL 2 2 6 100
TwoLayer_smallVoc 2 2 20 5
OneLayer 1 1 8 100
OneLayer_LARGE 1 1 32 100
Table 2: Transformer configurations used in our experiments.
A.2 Training and Regularization
All models are trained with the AdamW optimizer using a fixed learning rate of $6\times 10^{-4}$, $(\beta_{1},\beta_{2})=(0.9,0.95)$, and a batch size of $8$. We train for a fixed number of epochs.
A central aspect of our training setup is relatively strong weight decay. We use AdamW with an $\ell_{2}$ penalty $\lambda=0.1$ applied to all parameters with at least two dimensions (i.e., the token embedding matrix, attention projection matrices, and feed-forward weights), while excluding all bias terms and LayerNorm scale parameters from weight decay. This decoupled weight decay acts as our main form of explicit regularization in addition to dropout ($p=0.1$ in the attention, feed-forward, and embedding layers) and the small model sizes described in Section A.1. Together, these choices constrain effective capacity and discourage simple memorization of grammar-generated strings.
To further stabilize optimization, we apply gradient norm clipping with a maximum global norm of $1.0$ at every step. We do not use any learning-rate scheduling or warmup; the learning rate remains constant throughout training. Checkpoints are saved periodically and at the beginning and end of training, allowing us to analyze learning dynamics across epochs.
Appendix B Additional Proofs and Theorems
Proof of Proposition 3.9.
$\mathcal{L}(\theta)$ $=\sum_{x\in\Sigma^{*}}P(x)(-\log Q_{\theta}(x))$ (1)
$=\sum_{x\in\Sigma^{*}}P(x)(\log P(x)-\log P(x)-\log Q_{\theta}(x))$ (2)
$=\sum_{x\in\Sigma^{*}}P(x)\log\frac{P(x)}{Q_{\theta}(x)}-\sum_{x\in\Sigma^{*}}P(x)\log P(x)$ (3)
$=D_{\mathrm{KL}}(P~\|~Q_{\theta})-H(P)$ (4)
Proof of Theorem 4.1.
The decomposition can be constructed recursively. Given CFG $G=(\Sigma,\mathcal{N},\mathcal{S},\mathcal{P},\mathcal{W})$, the root node of the DAG – initially labeled with only $S$ – represents the entire grammar. If $S$ can generate itself through successive applications of rules of $G$, we add a self-loop from $S$ to itself.
Let $X\subseteq N$ be the subset of non-terminals on the right-hand side of any rule $S\rightarrow\alpha$. For each $A\in N$, let $G_{A}$ be the inner subgrammar generated by taking the closure of $A$ in $\mathcal{P}$ – that is, all the expansions $A\rightarrow\alpha$, all expansions of those non-terminals on the right-hand side of those rules, and so on. In the case that the result subgrammar is all of $G$, we can add $A$ as an additional label to the root node. Otherwise, $G_{A}$ is a proper inner subgrammar, in which case we assign it a node as a child of $S$. Inductively, this procedure is applied to each new subgrammar node (which by construction has strictly fewer non-terminals than its supergrammar). ∎
Proof of Theorem 4.3.
Theorems 4.3 and Corollary 4.4 are equivalent, for simplicity we directly prove Corollary 4.4.
Let $A_{1},\ldots,A_{k}$ be the top-level subgrammars of $G$. Let $S\rightarrow A_{i_{1,1},\ldots,i_{1,l_{1}}}$, …, $S\rightarrow A_{i_{r,1},\ldots,i_{r,l_{r}}}$ be all rules expanding $S$ in $\mathcal{P}$, with probabilities $p_{1},\ldots,p_{r}$ respectively. As we are directly proving Corollary 4.4, we assume $S$ expands only to non-terminals (by which we will also denote the top-level subgrammars; note that some of these may be $S$ itself if they are not proper subgrammars).
Denoting $P=P_{G}$ and $Q=Q_{\theta}$,
$D_{\mathrm{KL}}(P~\|~Q)$ $=\sum_{s\in\Sigma^{*}}P(s)\log\frac{P(s)}{Q(s)}$ (5)
$=\sum_{j=1}^{r}p_{j}\sum_{a_{j,1},\ldots,a_{j,l_{j}}}P(a_{j,1}\cdots a_{j,l_{j}})\log\frac{P(a_{j,1}\cdots a_{j,l_{j}})}{Q(a_{j,1}\cdots a_{j,l_{j}})}$ (6)
$=\sum_{j=1}^{r}p_{j}\sum_{a_{j,1},\ldots,a_{j,l_{j}}}P_{A_{j,1}}(a_{j,1})\cdots P_{A_{j,l_{j}}}(a_{j,l_{j}})\sum_{i=1}^{l_{j}}\log\frac{P_{A_{j,i}}(a_{j,i})}{Q(a_{j,i}|a_{j,1}\cdots a_{j,i-1})}$ (7)
$=\Big[\sum_{j=1}^{r}p_{j}\sum_{i=1}^{l_{j}}\sum_{a_{j,1},\ldots,a_{j,i-1}}P_{A_{j,1}}(a_{j,1})\cdots P_{A_{j,i-1}}(a_{j,i-1})\Big]\sum_{a}P_{A_{j,i}}(a)\log\frac{P_{A_{j,i}}(a)}{Q(a|a_{j,1}\cdots a_{j,i-1})}$ (8)
$=\sum_{i=1}^{k}\sum_{s}P_{G}(s|\epsilon)P_{G}(A|s)\sum_{a}P_{A_{i}}\log\frac{P_{A_{i}}(a)}{Q(a|s)}$ (9)
Corollary B.1.
Suppose $G$ has subgrammars $Z_{1},\ldots,Z_{l}$ as irreducible “leaf" subgrammars in its DAG subgrammar decomposition, and all rules evaluate to strings of only non-terminals, or only-terminals. Then
$D_{\mathrm{KL}}(P_{G}~\|~Q_{\theta})=\sum_{i=1}^{l}D_{\mathrm{KL}}(P_{G}~\|~Q_{\theta})_{Z_{i}}$
Proof of Theorem 4.6.
Let $G$ be a PCFG with top-level, proper subgrammars $A_{1},\ldots,A_{k}$. Summing over the top-level rules (expansions of $S$), suppose $S$ maps to a rule with $i$ recursive $S$’s with probability $p_{i}$ ($\sum_{i=0}^{N}p_{i}=1$ for some $N<\infty$). Then, $\mathbb{E}[R]=\sum_{i=1}^{N}p_{i}\cdot i$. Then by Corollary 4.5 (treating both proper subgrammars and recursive $S$ as top-level subgrammars), we have
$D_{\mathrm{KL}}(P_{G}~\|~Q_{\theta})$ $=\sum_{i=1}^{k}D_{\mathrm{KL}}(P_{A_{i}}~\|~Q_{\theta}(A_{i}))+\sum_{i=1}^{N}p_{i}\cdot iD_{\mathrm{KL}}(P_{G}~\|~Q_{\theta})$ (10)
$=\sum_{i=1}^{k}D_{\mathrm{KL}}(P_{A_{i}}~\|~Q_{\theta}(A_{i}))+\mathbb{E}[R]D_{\mathrm{KL}}(P_{G}~\|~Q_{\theta})$ (11)
$\implies~D_{\mathrm{KL}}(P_{G}~\|~Q_{\theta})$ $=\frac{\sum_{i=1}^{k}D_{\mathrm{KL}}(P_{A_{i}}~\|~Q_{\theta}(A_{i}))}{1-\mathbb{E}[R]}$ (12)
Theorem B.2.
For $G$ with outer subgrammar $A$, let $\bar{A}$ be its complement. The KL-divergence splits as a weighted sum:
$D_{\mathrm{KL}}(P_{G}~\|~Q_{\theta})=P_{G}(A)D_{\mathrm{KL}}(P_{A}~\|~Q_{\theta}|_{A})+P_{G}(\bar{A})D_{\mathrm{KL}}(P_{G}|_{\bar{A}}~\|~Q_{\theta}|_{\bar{A}})+D_{\mathrm{KL}}(P_{G}^{*}~\|~Q_{\theta}^{*})$
Where $D^{*}$, for $D\in\{P_{G},Q_{\theta}\}$ is the 2 valued distribution of whether $D$ outputs a string in $A$ or $\bar{A}$, $P_{A}$ is the language from CFG $A$, and $D|_{B}$ indicates the marginal distrubution of $D$ over strings of $B\in\{A,\bar{A}\}$.
Proof.
Writing $P$ for $P_{G}$ and $Q$ for $Q_{\theta}$ for legibility,
$D_{\mathrm{KL}}(P~\|~Q)$ $=\sum_{s\in A}P(s)\log\frac{P(s)}{Q(s)}+\sum_{s\in\bar{A}}P(s)\log\frac{P(s)}{Q(s)}$ (13)
$=P(A)\sum_{s\in A}P_{A}(s)[\log P(A)+\log P|_{A}(s)-\log Q^{*}(A)-\log Q|_{A}(s)]$ (14)
$+P(\bar{A})\sum_{s\in\bar{A}}P|_{\bar{A}}(s)[\log P(\bar{A})+\log P|_{\bar{A}}(s)-\log Q^{*}(\bar{A})-\log Q|_{\bar{A}}(s)]$ (15)
From which the final decomposition follows quite immediately by rearranging terms. ∎
Appendix C Definition of Grammars used for Experiments
In this section we properly introduce the PCFGs used for running the experiments.
KL decomposition example 1
[…]
KL decomposition example 2
[…]
Deeper Recursion
[…]
Outer Subgrammar Example
[…]
where the rules that are used for the unified subgrammar are highlighted in bold.
ABC Grammar
[…]
Nested Parentheses
[…]
Appendix D Additional Experimental results
D.1 ChatGPT-5 Instant Arithmetic Stress Test
We generate arithmetic expressions using integers uniformly sampled from 0–9 and the operators {+, -, *, / } are generated. Expression depth is defined as the maximum level of nested parentheses. Non-deep chains consist of 50 expressions of depth at most 2, concatenated by addition. Deep chains consist of single expressions with recursive nesting up to depth 7. Below are an example each:
Non-deep arithmetic expression:
((44)(19))+((6/3)(5/1))+((28)(8/5))+((5/9)(77))+((74)+(8/7))+((96)+(10))+((0/1)+(99))+((4/1)+(0+5))+((6+6)/(2/5))+((4/5)+(02))+((31)+(5+3))+((10)(76))+((25)(5/3))+((6+9)(6/1))+((1+4)/(6+9))+((9/7)(6+2))+((67)/9)+((4+1)+(73))+((53)(13))+((5+6)+4)+((52)+(00))+((67)8)+((5/2)+(4+6))+((5/5)(9/6))+((43)(87))+((7/3)(9+3))+((70)+(5/9))+((6/8)(2+0))+((0+6)/4)+((95)(39))+((0+1)+(94))+((77)(18))+((71)+9)+((40)+(08))+((6/9)(22))+((56)(8/4))+((35)/(4+2))+((34)(5+2))+((71)+(8/8))+((40)(9+7))+((3/6)(4/3))+((02)(1/9))+((08)(80))+((0/1)(2/8))+((9+5)(8/3))+((1+8)/(49))+((06)(2+4))+((5/6)+(2+0))+((27)(2/2))+((8+8)2)((4*4)*(1-9))+((6/3)*(5/1))+((2-8)*(8/5))+((5/9)*(7*7))+((7-4)+(8/7))+((9-6)+(1-0))+((0/1)+(9-9))+((4/1)+(0+5))+((6+6)/(2/5))+((4/5)+(0-2))+((3*1)+(5+3))+((1-0)-(7-6))+((2*5)*(5/3))+((6+9)-(6/1))+((1+4)/(6+9))+((9/7)-(6+2))+((6-7)/9)+((4+1)+(7-3))+((5-3)-(1*3))+((5+6)+4)+((5*2)+(0-0))+((6*7)*8)+((5/2)+(4+6))+((5/5)*(9/6))+((4-3)*(8*7))+((7/3)*(9+3))+((7-0)+(5/9))+((6/8)-(2+0))+((0+6)/4)+((9-5)-(3-9))+((0+1)+(9-4))+((7-7)*(1-8))+((7-1)+9)+((4-0)+(0*8))+((6/9)*(2-2))+((5-6)-(8/4))+((3*5)/(4+2))+((3*4)-(5+2))+((7-1)+(8/8))+((4*0)-(9+7))+((3/6)-(4/3))+((0-2)-(1/9))+((0-8)*(8*0))+((0/1)*(2/8))+((9+5)*(8/3))+((1+8)/(4-9))+((0*6)*(2+4))+((5/6)+(2+0))+((2*7)-(2/2))+((8+8)*2)
Result: $\frac{707449}{1260}$
Deep arithmetic expression:
(((((((3+8)+(51))((16)+(5+3)))+(((82)(38))+((29)(4+5))))((((1/7)(64))((7+3)(6+6)))(((83)(1+8))+((59)+(7/1)))))+(((((86)/(53))((80)(80)))+(((89)+(36))((9/8)/(78))))/((((74)(2+2))((35)/(92)))/(((6/8)+(55))((41)(8+8))))))((((((40)/(48))((80)(31)))+(((77)(4/7))((70)(0/7))))((((86)/(8+7))((8/8)+(8/4)))(((5+5)(98))((9/2)/(39)))))+(((((98)+(21))((4+3)/(95)))/(((22)(43))((66)(6+9))))((((8/8)(33))/((8+0)+(9/1)))(((21)6)((1+5)/8))))))(((((((3+8)+(5-1))-((1-6)+(5+3)))+(((8-2)-(3-8))+((2*9)*(4+5))))*((((1/7)-(6*4))*((7+3)*(6+6)))-(((8*3)*(1+8))+((5-9)+(7/1)))))+(((((8*6)/(5-3))*((8*0)-(8-0)))+(((8-9)+(3-6))-((9/8)/(7*8))))/((((7-4)*(2+2))-((3-5)/(9-2)))/(((6/8)+(5*5))*((4-1)-(8+8))))))-((((((4-0)/(4-8))*((8-0)-(3-1)))+(((7*7)*(4/7))*((7*0)-(0/7))))-((((8*6)/(8+7))-((8/8)+(8/4)))-(((5+5)*(9*8))-((9/2)/(3-9)))))+(((((9-8)+(2*1))-((4+3)/(9-5)))/(((2*2)*(4*3))-((6-6)-(6+9))))*((((8/8)-(3*3))/((8+0)+(9/1)))*(((2*1)*6)*((1+5)/8))))))
Result: $\frac{892410719}{448320600}$
D.2 Recursive Decomposition Experiments
Figure 4: Two-layer Transformer showing the impact of the probability of recursion.
D.3 Pretraining Results
This appendix provides the detailed results referenced in the main text. All experiments compare transformers trained from scratch against those pretrained on a subgrammar before continuing on the full grammar. Figure 5 shows that no matter which subgrammar is chosen, when later training on the full grammar, it is not forgotten.
(a) Pretraining on an infix subgrammar
(b) Pretraining on a suffix subgrammar
Figure 5: Examples of pretraining on differently placed subgrammars using ABC Grammar.
Figure 6 illustrates the distribution of KL-divergences across 30 seeds when training directly versus with 10 epochs of subgrammar pretraining. Pretraining consistently shifts the distribution toward lower KL.
(a) Two-layer Transformer
(b) Four-layer Transformer
Figure 6: Distribution of final KL value of pretraining versus training from scratch
Table 3 reports average cosine similarity across attention and MLP layers, on three types of test sequences: (i) sequences consisting solely of subgrammar subsequences, (ii) sequences with no subgrammar subsequences, and (iii) sequences mixing subgrammar and other subsequences.
Attention MLP
Sequences with subgrammar only
From Scratch 0.660 0.635
With Pretraining 0.743 0.611
Percentage change (%) +12.6 -3.9
Sequences without subgrammar
From Scratch 0.835 0.837
With Pretraining 0.876 0.841
Percentage change (%) +4.9 +0.5
Sequences with subgrammar
From Scratch 0.726 0.501
With Pretraining 0.687 0.543
Percentage change (%) -5.7 +8.4
Table 3: Average cosine similarity [-1, +1] across attention and MLP layers of a two-layer Transformer when pretraining for 10 epochs. Percentage change indicates the relative difference between models trained from scratch and with pretraining.
D.4 Generalization and Prefix Experiments
This appendix provides the detailed figures referenced in the main text. They compare how different valid prefixes (shallow vs. deeply recursive) and malformed prefixes affect model stability, showing that deeply recursive but valid prefixes can degrade performance even more than ungrammatical ones.
(a) Prefix = $(a)(a)(a)(a)(a)(a)$
(b) Prefix = $(((((((((a))))))))$
(c) Wrong prefix = $(a)(a)(a))(aa)(a)(a)$
Figure 7: Comparison of different prefixes for recursion type 2